# Reflected Brownian Motion in a wedge: sum-of-exponential absorption probability at the vertex and differential properties

### v1

@@ 4 Absorption probability via compensation approach
This section deals with the case where $\alpha\in\mathbb{N}^{*}$. The aim is to show that the absorption probability is a sum of exponentials and to calculate precisely all the coefficients of this sum. To that end we invert the Laplace transforms and we explain the recursive compensation phenomenon which appears in this sum.
@@
When $\alpha\in\mathbb{N}^{*}$ and the simple root condition (38) holds, we invert the Laplace transforms $\varphi_{1}$ and $\varphi_{2}$ obtained in Theorem 9 by performing a partial fraction decomposition. Therefore, remembering that $\varphi_{1}$ and $\varphi_{2}$ are defined in (12) as the Laplace transforms of the escape probability on the boundaries, the absorption probabilities starting from the boundaries can be written as sum-of-exponential and are explicitly given by
@@
[…]
@@
First we assume that $\alpha=n$ and for all $j\in\{1,\dots,2\alpha-2\}$, $\theta-2\delta+j\beta\not\equiv 0\text{ mod}(\pi)$. Theorem 9 gives an explicit expression of the Laplace transform $\varphi$. Performing a partial fraction decomposition of $\varphi(x,y)=\frac{L(x,y)}{P(x)Q(y)}$ it is possible to invert the Laplace transform. Recall the definition of $P$ and $Q$ given in (31) and remark that $\{s_{2}\bm{\mathrm{q}}^{i}:i=0,\dots,\alpha-1\}=\{s_{1}/\bm{\mathrm{q}}^{i}:i=1,\dots,\alpha\}$ since $s_{1}=s_{2}\mathrm{q}^{\alpha}$, we obtain $\alpha^{2}$ constants $\tilde{c}_{i,j}$ such that
@@
We assume $\alpha\in\mathbb{N}^{*}$ and condition (38). We recall that the constants $a_{k}$, $b_{k}$ are defined in (46). The constants $c_{k}$ introduced in Theorem 12 are determined by the recurrence relations
@@
In this paragraph, we explain the compensation phenomenon. By using an analytic approach, we showed in Theorem 12 that when $\alpha\in\mathbb{N}^{*}$ and (38) holds the absorption probability is
@@
This last paragraph deals with the case where $P$ or $Q$ have double roots, i.e. when for some integer $j\in\{1,\dots,2\alpha-2\}$, $\theta-2\delta+j\beta\equiv 0\text{ mod}(\pi)$, see Lemma 8. The number of cases to handle to give a general explicit formula is too big. Nonetheless, we can give the general shape of the absorption probability: if $\alpha\in\mathbb{N}^{*}$, the absorption probability can be written as
@@ Acknowledgments
This project has received funding from Agence Nationale de la Recherche, ANR JCJC programme under the Grant Agreement ANR-22-CE40-0002.

### v2

@@ 4 Absorption probability via compensation approach
This section deals with the case where $\alpha\in\mathbb{N}$. The aim is to show that the absorption probability is a sum of exponentials and to calculate precisely all the coefficients of this sum. To that end we invert the Laplace transforms and we explain the recursive compensation phenomenon which appears in this sum.
@@
When $\alpha\in\mathbb{N}$ and the simple root condition (41) holds, we invert the Laplace transforms $\varphi_{1}$ and $\varphi_{2}$ obtained in Theorem 10 by performing a partial fraction decomposition. Therefore, remembering that $\varphi_{1}$ and $\varphi_{2}$ are defined in (12) as the Laplace transforms of the escape probability on the boundaries, the absorption probabilities starting from the boundaries can be written as sum-of-exponential and are explicitly given by
@@
$(a_{2k},b_{2k}):=\left(\bm{\mathrm{x}}\left(s_{1}/\bm{\mathrm{q}}^{k}\right),\bm{\mathrm{y}}\left(s_{1}/\bm{\mathrm{q}}^{k}\right)\right)\quad\text{and}\quad(a_{2k+1},b_{2k+1}):=\left(\bm{\mathrm{x}}\left(s_{1}/\bm{\mathrm{q}}^{k+1}\right),\bm{\mathrm{y}}\left(s_{1}/\bm{\mathrm{q}}^{k}\right)\right)$ (50)
@@
First we assume that $\alpha=n$ and for all $j\in\{1,\dots,2\alpha-2\}$, $\theta-2\delta+j\beta\not\equiv 0\text{ mod}(\pi)$. Theorem 10 gives an explicit expression of the Laplace transform: one can write $\varphi(x,y)=\frac{L(x,y)}{P(x)Q(y)}$ where $L$ is a polynomial given by (42) and perform a partial fraction decomposition of $\varphi$. It is then possible to invert the Laplace transform. Recall the definition of $P$ and $Q$ given in (34) and remark that $\{s_{2}\bm{\mathrm{q}}^{i}:i=0,\dots,\alpha-1\}=\{s_{1}/\bm{\mathrm{q}}^{i}:i=1,\dots,\alpha\}$ since $s_{1}=s_{2}\mathrm{q}^{\alpha}$, we obtain $\alpha^{2}$ constants $\tilde{c}_{i,j}$ such that
@@
We assume $\alpha\in\mathbb{N}$ and condition (41). We recall that the constants $a_{k}$, $b_{k}$ are defined in (50). The constants $c_{k}$ introduced in Theorem 13 are determined by the recurrence relations
@@
In this paragraph, we explain the compensation phenomenon. By using an analytic approach, we showed in Theorem 13 that when $\alpha\in\mathbb{N}$ and (41) holds the absorption probability is
@@
This last paragraph deals with the case where $P$ or $Q$ have double roots, i.e. when for some integer $j\in\{1,\dots,2\alpha-2\}$, $\theta-2\delta+j\beta\equiv 0\text{ mod}(\pi)$, see Lemma 9. The number of cases to handle to give a general explicit formula is too big. Nonetheless, we can give the general shape of the absorption probability: if $\alpha\in\mathbb{N}$, the absorption probability can be written as
@@ Acknowledgments
This project has received funding from Agence Nationale de la Recherche, ANR JCJC programme under the Grant Agreement ANR-22-CE40-0002. The authors would like to thank the anonymous referee for his/her helpful comment on the irreducibility of the kernel.